June 2004

Sunday Monday Tuesday Wednesday Thursday Friday Saturday
 
 
 
 
 
 
1
Last 2 Spectrometer planes installed - 122 & 121. Plane 121 has cables attached and light-leak checked by 2pm.

Near Hall LCW cooling system turned on - no water leaks in racks. Special filters installed to clean up the system.

Complete installation of top level of concrete blocks around DK pipe. Deliver first 4 Blu Blocks to Absorber Enclosure (88 total).
First 2 Blu Blks set in place.

Sump #2 fails again.
2
Ongoing Near Det electronics installation and shake-down of RPS system.

5 Blu Blks delivered to Absorber Enclosure. Continue Blu Blk installation.
3
Ongoing Near Det electronics installation and shake-down of RPS system.

Delivered 4 Blu Blks to Ab Encl. Continue installation of blocks.

Big electric fork truck returns from vendor, with an upgraded motor. Carries one Blu Blk up the hill OK.  During 2nd trip up hill with 2nd Blu Blk, the new motor overheats. Transfer Blk to winch cart, and cease using fork truck to carry loads up the hill.  For now, use it to transfer Blu Blks from cart at top of hill into Ab Encl.
4
Ongoing Near Det electronics installation and shake-down of RPS system.

5 Blu Blks delivered to Ab Encl. Continue installation of blocks.

Remove Sump #2 for inspection & repair.
5

 
 
 
 
6
Projected Work:
• Complete Spectrometer Installation 6/1. Then halt detector installation, for meeting at Week-in-the-Woods and tech staff catch-up work at New Muon.
• Start to bring up Near Hall Water Cooling system - for racks and magnet coil.
• Continue installing Absorber.
7
Ongoing Near Det electronics installation and shake-down of RPS system.

Delivered 3 Blu Blks to Ab Encl.  Continue installation of blocks.
8
Ongoing Near Det electronics installation and shake-down of RPS system.

Delivered 5 Blu Blks, and the 2 halves of the steel pan for the 2ndry containment system (for holding activated water from leaks in Absorber Core cooling system).

Completed installation of the 1st two layers of Blu Blks - 24 blocks in 2 layers of 12 each. 2ndry pan sits on 2nd layer.

Start grouting outer edges of installed Blu Blks (to assist in air containment).
9
Electricians connecting ground strap between ND Rack frames and ground buss.  (this is separate from the electronics backplane ground).

Continue grouting of 1st 2 layers of Blu Blks. Also shim and grout surface to level, so that 2ndry containment pan sits level.

6 Blu Blks lowered, to be staged in the enclosure.

Another test of big Fork Truck, now with additional fans to help keep motor cool. This does seem to help, but only to a point.
10
Plane 120 is lowered and staged onto the cart, but not installed on the detector. Plan for Cal Plane Inst is to always have the 1st plane of the day staged underground at the end of the previous workday.

Lowered 5 more Blu Blks to be staged.

Continue shim and grount for leveling of 2ndry pan.

Use Big Fork Truck to deliver a Blu Blk, then give it 20min rest, with fans blowing, to cool off before delivering the next - as reccommended by Service Rep.  But on delivery of 2nd Blk, the motor still seriously overheats.
11
Planes 119 and 118 are brought to MSB and laid on the floor (on strongbacks). Plane 120 is installed on the detector. Planes 119 waits on the cart until 120 is cabled and checked on Monday.

Complete inst of rack frame grounding. Begin inst of ground lines and buss for electronics crates.

Lowered 2 Blu Blks.  Now have 14 Blks staged underground.

Completed grout and level of 1st 2 layers of Blu Blks.
12

 
 
 
 
13
Projected Work:
• Begin Calorimeter Plane installation, at 3 Planes per day, beginning at 7am daily.
• Continue Detector Commissioning, begining at 8am daily - every plane now needs cabling and checkout before the next can be installed.
• Continue Absorber Installation.
14
3 Planes installed, cabled and light-leak checked: 120, 119, 118.

Continue inst of electronics signal grounding

4 Blu Blks staged. Install 2ndry containment pan. Start layer 3 - set 4 Blu Blks down the center of the pan.
15
3 Planes installed, cabled and light-leak checked: 117, 116, 115.

Complete inst of electronics signal grounding

4 Blu Blks staged. Continue layer 3 - set steel plate along the sides of the pan.
16
3 Planes installed, cabled and light-leak checked: 114, 113, 112.

2 Blu Blks staged. Core carrier plates delivered to Enclosure. Complete layer 3 Blu Blks. Install survey laser for leveling Blu Blks. Mark location for Hadron Monitor support rails.
17
3 Planes installed, cabled and light-leak checked: 111, 110, 109.

Safety inspection of electronics racks results in Operational Readiness Clearance, which means the racks can be left powered, unattended.

Install Hadron Monitor support rails, with survey location check. Start layer 4 Blu Blks.

Removed big forktruck & battery packs - it will no longer be used for installation.
18
3 Planes installed, cabled and light-leak checked: 108, 107, 106.

Complete layer 4 - there are now 48 Blu Blks set in place. Core carrier steel plates set in place.
19

 
 
 
 
20
Projected Work:
• Continue 3 Plane per day Installation, beginning at 7am daily.
• Continue Detector Commissioning, begining at 8am.
• Continue Absorner Installation
21
2 Planes installed, cabled and light-leak checked: 105, 104.

Shaft crane makes burning odor shortly after Plane 104 is delivered underground. Arrival of repair, and repair work, takes until 3:30pm.

1 Blu Blk lowered while crane repair crew is present. Grout layer 4 blks. Start hand-stack of bricks between upst wall and concrete shield blocks.
22
3 Planes installed, cabled and light-leak checked: 103, 102, 101.

4 Blu Blks staged underground. Continue hand stack. Start inst of sliding door over Hadron access slot.
23
3 Planes installed, cabled and light-leak checked: 100, 99, 98.

3 Blu Blks staged underground - for a current total of 8. But had to remove 2 core carrier plates and 4 Blu Blks from layer 4 to be able to install Air Isolation Box.

Electricians working on Absorber area Interlocks, and also bundling HV cable for Muon detectors.
24
3 Planes installed, cabled and light-leak checked: 97, 96, 95.

Waiting for Air Isolation Box before continuing with absorber steel layers. Start to install lower level of outer B concrete blocks on bothe east and west sides. 4 B blocks lowered and set.

At 2:30pm, after most shaft crane operations for the day were complete, the shaft crane controls exhibit problems. CranePro called and will be here in the AM.
25
CranePro arrives at 8am and observes operations of shaft crane until 1pm - no repairs found to be necessary. Shaft crane deliveries continue while the crane is being observed.

3 Planes installed, cabled and light-leak checked: 94, 93, 92.

Moved 4 B blocks to Absorber Enclosure, installed along sides of Pile.

Hard Rock has been coring a hole between the sump and the deisel pump located just upstream of the shaft, for use of the pump line in emergencies when all the sump pumps fail.

26

 
 
 
 
27
Projected Work:
• Continue 3 Plane per day Installation, beginning at 7am daily.
• Continue Detector Commissioning, begining at 8am.
• Continue Absorber Installation.
28
3 Planes installed, cabled and light-leak checked: 91, 90, 89.

2 B blocks staged underground. Air isolation box delivered to enclosure and installed. Blu Blks and Core Plates which were removed to allow Box installation are replaced.

Electricians begin pulling HV cable for Muon Alcoves.
29
3 Planes installed, cabled and light-leak checked: 88, 87, 86.

Delivered 1st half of Aluminum Core to Absorber Enclosure. 2nd half, and the Steel Core portions, are in MSB.  Grouted the Blu Blks and core carrier plates installed yesterday. Lower row of B Blocks along both sides are in place, but not yet shimmed level.
30
3 Planes installed, cabled and light-leak checked: 85, 84, 83.

Installed 1st Al core section, then lowered, delivered and installed both the 2nd Al core sectionand the 1st steel core section.

Electricians complete pulling cables between Muon Alcoves and Muon readout racks.

Hard Rock completes coring of a trench for the emergency backup diesel sump pump.